• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sarty

    Log entry time 23:55:01 on July 24, 2006

    Entry number 175160

    keyword=Shift summary - swing

    16:00 Arrive with MCC tuning beam to come back / recover from the past day's down resulting from vacuum problem in North Linac.

    17:00-18:00 Take a controlled access with Alex to reboot ROC3 (HRS-L DAQ system) - see separate Halog entry. MCC continued to tune.

    22:05 While working to ensure HRS-L is set to proper momentum, we had some problems (see separate halog entries this shift) - then had HRS-L Q2 Quench fault. Scot came in, and had an access to fix. During access, Alex also hooked up T9 triggers (per Bob M's request - thanks), and also fixed a loose cable connection to one of the VDC HV supply current readouts.

    By end of shift...still waiting for beam. MCC has been working on trying to get Hall C's beam acceptable, and not having great luck apparently. The plan was for them to stop trying for better tune to Hall C tonight at midnite (this shift's end) so that Hall C can use whatever they can get; then our beamline will start to be tuned.

    Good luck to next shift :)